Can you please help me find each measure using the given measures from triangle ABC. I need to round the angle measures to the nearest degree and side measures to the nearest tenth for this problem: IF b=14, c=12, and the measure of angle A=62, find a.
Answer by stanbon(48535)

Can you please help me find each measure using the given measures from triangle ABC. I need to round the angle measures to the nearest degree and side measures to the nearest tenth for this problem: IF b=14, c=12, and the measure of angle A=62, find a.
----------
a^2=b^2+c^2-2bc(Cos(A))
a^2 = 14^2 + 12^2 -2*14*12(Cos(62))
a^2 = 182.25755...
a= 13.50
========
Cheers,
Stan H.